OpenStudy (piercetheveil47):

There are (7^13^)3 ⋅ 7^0 strawberries in a field. What is the total number of strawberries in the field? 1 7^16 7^17 7^39

OpenStudy (piercetheveil47):

*(7^13)^3

OpenStudy (mathstudent55):

What is 7^0 = ?

OpenStudy (piercetheveil47):

0?

OpenStudy (piercetheveil47):

@mathstudent55

OpenStudy (piercetheveil47):

jk 1

OpenStudy (cwrw238):

(7^13)^3 = 7 ^ (13*3)

OpenStudy (mathstudent55):

No. A number to the zero power equals 1, except 0^0 which is undefined. 7^0 = 1

OpenStudy (mathstudent55):

Also, when you raise a power to a power, you multiply powers.

OpenStudy (mathstudent55):

Example: